OpenStudy (anonymous):

If (a+b)=(6/7) and (a/c)=(2/5) what is the value of 3b+c in terms of a ? a) 12a b) 9a c) 8a d) 6a e) 4a

OpenStudy (saifoo.khan):

(a+b)=(6/7) -equation 1 Solve for a. a = 6/7 - b Now use this value and insert in equation 2 (a/c)=(2/5) \[\Large \frac{ \frac 67 - b}{c} = \frac 25\] Now solve for c.

OpenStudy (anonymous):

how do i solve for a

OpenStudy (saifoo.khan):

-Make (a) the subject

OpenStudy (saifoo.khan):

If i'm not wrong it's a SAT question?

OpenStudy (anonymous):

Yes its a SAT question

OpenStudy (saifoo.khan):

Dang! i'm going the long way. -_- Let me show you again.. easier method..

OpenStudy (saifoo.khan):